Write the standard form of the equation of the line through the pair of points (-4,-8) (-8,3). Write the equation using only integer coefficients
Answer by mananth(16949)

x1 y1 x2 y2
-4 -8 -8 3

slope m = (y2-y1)/(x2-x1)
( 3 - -8 )/( -8 - -4 )
( 11 / -4 )
m= -11/ 4

Plug value of the slope and point ( -4,-8 ) in
Y = m x + b
-8.00 = 11 + b
b= -8 - -11
b= 3
So the equation will be
Y = -11/4 x + 3
multiply by 4
4y=-11x+12
Now you can write in any form you wish
m.ananth@hotmail.ca